atoms of Carbon dioxide gas decay from upper level to lower level with a spread in frequency is about 10kHz. calculate the line shape if the broadening mechanism is found Doppler broadening?
1
Expert's answer
2020-05-21T13:36:44-0400

The line shape is Gaussian with the standard deviation of


"\sigma=\sqrt{\frac{kT}{Mc^2}}\omega_0."


The distribution of frequencies will be


"L=\frac{1}{\sigma\sqrt{2\pi}}\text{ exp}\bigg[-\frac{(f-f_0)^2}{2\sigma^2}\bigg]."